In 1999, 18.2 million households had a dvd player. If this populaion grows at 25% annually, how many households will have a dvd player in 2004 (5 years later)? A= P(1+r)^t
I got: 55.54 million. Is this correct?
Thanks~

Answer by stanbon(75887)

In 1999, 18.2 million households had a dvd player. If this populaion grows at 25% annually, how many households will have a dvd player in 2004 (5 years later)? A= P(1+r)^t
----------
A(5) = 18.2(1+0.25)^5
A(5)= 18.2(3.05175...)
A(5)= 55.54199... million
==========
Cheers,
Stan H.